U N I V E R S A L C A L E N D A R Author Valentin Hristov (Bulgaria)

Year in the century

Instructions 00 01 02 03 04 05

The date Thursday

for use 06 07 08 09 10 11

October 4, 1582 (old style)

1.

12 13 14 15 16

is followed by the date Friday

Year in century 17 18 19 20 21 22

October 15, 1582 (new style).

(intersect!) ↓ 23  24 25 26 27

Month → K 28 29 30 31 32 33

In Bulgaria the date Thursday

K - key day 34 35 36 37 38 39

March 31, 1916 (old style)

(remember K!)

40 41 42 43 44

is followed by the date Friday

2.

45 46 47 48 49 50

April 14, 1916 (new style).

K ← Century 51  52 53 54 55

(remember 56 57 58 59 60 61

The leap years and

col. with K!)

62 63 64 65 66 67

exact centuries

3. 68 69 70 71 72 are in bold.

D ← Date 73 74 75 76 77 78

January and February

D - day 79 80 81 82 83

in leap years are also

(in remembered 84 85 86 87 88 89

in bold – Jan and Feb.

column)

90 91 92 93 94 95

Full centuries

Month 96 97 98 99 Old style New style
Jun  Mo Tu We Th Fr Sa Su 1 8 15 22 29
Dec Sep Tu We Th Fr Sa Su Mo 2 9 16 23 30 18 22
Jul Apr Jan We Th Fr Sa Su Mo Tu 3 10 17 24 31
Oct  Jan Th Fr Sa Su Mo Tu We 4 11 18 25  15 19 23
May  Fr Sa Su Mo Tu We Th 5 12 19 26  16 20 24
Aug  Feb Sa Su Mo Tu We Th Fr 6 13 20 27
Nov  Mar Feb Su Mo Tu We Th Fr Sa 7 14 21 28  17 21 25
Day  Date

Detailed description of the steps

All steps use the central rectangle with the names of the days of the week.

1. Intersect the horizontal row containing the month (left) and the vertical column containing the last two digits of the year (above) to find the key day K. Remember it just as a symbol to be used in Step 2! (Be careful if the month is January or February. For leap years (see Note: below) use Jan and Feb.)

2. Find the full centuries (right) in one of the rectangles in accordance with the style – old or new. In the horizontal row with the centuries find the key day K from Step 1. Remember the column containing K in the central rectangle for Step 3!

3. The rectangle, used in Step 2 for old style centuries, is now used for the dates of the month. Together with the column remembered in Step 2 it gives the month calendar, i.e. intersect the horizontal row with the date of the month and the remembered column to find the day of the week.

Note: Leap years in the old as well as in the new style are those which can be divided by 4 without remainder. There are exceptions only for the new style: the exact centuries (ending with 00 ), which cannot be divided by 400 without remainder, are not leap years.